OpenStudy (anonymous):

4. Part A: Explain why the x-coordinates of the points where the graphs of the equations y=2-x and y = 4x + 3 intersect are the solutions of the equation 2-x = 4x + 3 Hint: Think about graphing (5 points) Part B: Make tables to find the solution to 2-x = 4x + 3. Take the values of x between −3 and 3. (5 points) @Sadworld

OpenStudy (anonymous):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 2-x and y = 4x + 3 intersect are the solutions of the equation 2-x = 4x + 3. Hint: Think about graphing (5 points) Part B: Make tables to find the solution to 2-x = 4x + 3. Take the values of x between −3 and 3. Be sure to state the solution. (5 points) y = 2-x y = 4x + 3 x y 1 0 -1 -2 x y 1 0 -1 -2

OpenStudy (anonymous):

I don't know to answer Part A asides from that both equation are part of a system. Use substitute x in the equation 2-x = 4x + 3 to find the y of all x.

OpenStudy (anonymous):

can u explain

OpenStudy (anonymous):

Okay, give me a min to type.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Do you think you can add a print screen? I want to be sure

OpenStudy (anonymous):

what is that

OpenStudy (anonymous):



OpenStudy (anonymous):

what about b

OpenStudy (anonymous):

Just solve both independently to get the value of y.

OpenStudy (anonymous):

explain please

OpenStudy (anonymous):